Status Report

NASA Space Station On-Orbit Status 15 July 2004

By SpaceRef Editor
July 17, 2004
Filed under , , ,
NASA Space Station On-Orbit Status 15 July 2004

SpaceRef note: This NASA Headquarters internal status report, as presented here, contains additional, original material produced by (copyright © 2004) to enhance access to related status reports and NASA activities.

All ISS systems continue to function nominally except those noted previously or below.  

Update on Russian Segment (RS) software upgrading:   The attempted uploading of the new SM 7.02 software (s/w) to the Service Module (SM) computers this morning failed for currently unknown reasons and will be reattempted tomorrow after careful analysis by TsUP/Moscow.  Onboard systems continue to run on the existing 7.01 s/w.   [After ground and crew deactivated the Vozdukh CO2 scrubber, Elektron O2 generator and BMP impurities filtration unit starting at 4:50am EDT (SKV-1 air conditioner was already off), next steps were to be the actual s/w upload, starting with the #3 subset (lane) of the TVM (Terminal Computer), later of the TsVM (Central Computer), followed by restart of each computer and subsequent loading of the remaining lanes, with another restart (i.e., three restarts total).  However, the TVM sub computer “refused to accept” the new s/w, even after two tries, and activities are currently on hold.  Deactivated ECLS systems were restored to operation (except for the SKV-1 air conditioner, see below).]

Update on SM cooling pump failure:   More bad luck in the RS: Yesterday a pump in the SM’s no. 2 internal cooling loop (KOB-2) unexpectedly suffered a hard failure, causing the associated no. 1 pump panel (4SPN1) to transition to the #2 pump panel (3SPN2) of KOB-1.  Fruitless TsUP attempts to command a transition back to 4SPN1 showed that three of the four pump panels (two per internal loop) are not operational at this time, rendering the SM’s internal cooling system zero-fault tolerant (i.e., one more failure = no cooling).  Removal & replacement (R&R) of the failed pump panel will be attempted tomorrow, followed by switching back to automatic mode from manual mode.  As an impact of the KOB-2 pump failure, the SKV-1 air conditioner is off (at this time, there is no issue with the humidity levels).    [Each of the four pump panels (27 kg each) has two pumps hard welded to it, and the automatic algorithm (logic) in charge does not allow a panel to run on only one of them, thus the refusal to transition back.  Spares are on board, but the replacement panel is a new lighter-weight design (its pumps, weighing 10 kg each, are not hard welded to the base but individually replaceable without requiring swap-out of the entire assembly), but that requires new procedures for the R&R to be developed ASAP and uplinked to the crew overnight.]

Update on US EMU spacesuit troubleshooting:  As a small change in plan for next Monday’s (7/19) scheduled IFM (inflight maintenance) due to the late decision to maintain integrity of suit 3005 for as long as possible, the troubleshooting will be done only on EMU 3013 and will not involve loosening of the screws of the water pump housing.  Instead, the housing will actually be opened and the pump removed for inspection, documentation, analysis and determination of subsequent work.  This is a “two-handed” task that can be done by Mike Fincke alone.

Expedition 9 Flight Crew Plans can be found at

First thing in the morning, before breakfast and exercise, CDR Padalka completed another session with the periodic Russian MedOps test “Hematokrit” (MO-10), measuring red blood cell count of the blood.   [The blood sample was drawn from a finger with a perforator lancet, then centrifuged in two microcapillary tubes in the M-1100 kit’s minicentrifuge, and its hematocrit value was read off the tubes with a magnifying glass.  It is a well-known phenomenon of space flight that red blood cell count (normal range: 30-45%) tends to go down over time.  After the exam, the data were saved in the IFEP software (In-Flight Examination Program) on the MEC (Medical Equipment Computer).]

Afterwards, still going hungry, Padalka also took the MBI-1 SPRUT-K test, part of Russian medical research on the distribution and behavior of human body fluids in zero gravity.  Supported by Laptop 3, the data were recorded on “Profilaktika” memory cards, along with this morning’s hematocrit data and yesterday’s body mass values.  Afterwards, LT3 was powered down.  [Experiment requisites are the Sprut (“squid”) securing harness, skin electrodes (cuffs), and the payload computer for control and data storage.  The “Penguin” suit or “Braslet-M cuffs, if worn, have to be taken off first.  Electrode measurements are recorded at complete rest and relaxed body position.  Assistance from the FE was not required.]

Previous Reports

ISS On-orbit Status [HQ]
ISS Status [JSC]
Shuttle Processing [KSC]

Russian biomed assessments on the CDR continued later in the day, when he conducted his second session with the cardio experiment PZEh MO-1 (Study of the Bioelectric Activity of the Heart at Rest), with Fincke assisting as CMO.   [During the 30-min. test, the crew tagged up with ground specialists on a Russian ground site (RGS) pass on Daily Orbit 2 and downlinked data from the Gamma-1M ECG (electrocardiograph) for about 5-6 minutes.]

FE/SO Mike Fincke continued his current assignment to collect and pack US cargo items designated for disposal on Progress M-249/14P.  Later, he and Padalka started trash stowing in the Progress, guided by the IMS (inventory management system).

Gennady also tagged up with stowage specialists at TsUP via S-band on his weekly IMS teleconference to discuss open issues concerning identification of equipment and storage locations for updating the IMS database.   [Today’s exchange, at 5:00am EDT, focused on identification of stowage locations associated with Padalka’s installation of new enclosures behind FGB panels on 7/12 and various excessed equipment.]

The CDR continued the current round of monthly preventive maintenance on Russian segment (RS) air ventilation systems, today working an hour in the Funktsionalnyi-Grusovoi Blok (FGB) for another cleanup of the fan screens of the three SOTR (thermal control system) gas-liquid heat exchangers (GZhT).

The ISS-9 crew performed the mandatory OBT emergency egress drill for the case of rapid cabin depressurization.   [Objectives of this Russian-led exercise are to (a) exercise and review depress response procedures, (b) practice crew coordination, (c) review questions and crew comments on the FDF (flight data file) and ISS hardware, and (d) ensure familiarity with emergency books storage location.  The exercise, which followed a scripted scenario and did not require actual manipulation of hardware, took about 1.5 hrs.  The drill started with the crew being informed about the initial pressure and the current pressure values in the station, the depress rate, and other data required to “perform” the procedures as per FDF EMER-1 book.  During the session, the crew simulated executing the planned emergency procedures while moving about the station.  For the case of an onboard fire and for emergency descent, there are other mandatory emergency drill OBTs.]

At the Lab RWS (robotics work station), Fincke re-connected the UOP-DCP (utility outlet panel-to-display & control panel) bypass power cable for the MSS (mobile service system) for the subsequent use of the SSRMS (Space Station Remote Manipulator System) video system.  Later tonight, the RWS and SSRMS will be powered up from the ground for an MSS camera characterization activity during crew sleep.   [The characterization essentially involves a settings calibration of camera eye points, a focus-iris-zoom test and a checkout of camera lights.]

Mike conducted the routine maintenance of the SM’s life support system (SOZh), while Gennady prepared the daily IMS “delta” file update, which had been added to the discretionary Russian task list.

The crew completed their daily 2.5-hr. physical exercise program on TVIS treadmill, RED expander and VELO ergometer with load trainer.

At 3:45pm, the crew is scheduled to engage in an interactive audio PAO ship-to-ship call with the crew of the ten-day NEEMO 6 mission (NASA Extreme Environment Mission Operations) currently underway in the “Aquarius” underwater research facility.  Video will be taken on both sides and later edited/”packaged” by PAO for NASA TV.   [Aquarius, similar in size to the ISS SM and owned by NOAA, is the only undersea research laboratory in the world.  The 45-ft long by 13-ft wide habitat sits on the ocean floor three miles off Key Largo, Florida, near deep coral reefs, 62 ft beneath the surface.  The new four-member NASA crew is headed by Native American Astronaut John Herrington; the other crewmembers are NASA engineer Tara Ruttley and Astronauts Nicholas Patrick and Douglas Wheelock.]

The lens change (to 180mm) and refocusing of the EarthKAM (EK) digital camera in the Node yesterday are reportedly yielding excellent image quality.  The current EK run has taken photo requests from 26 elementary, middle, and high schools in the US, Canada, Germany, Spain, and Japan. 

Previous NASA ISS On-orbit Status Reports can be found here. Previous NASA Space Station Status Reports can be found here. Previous NASA Space Shuttle Processing Status Reports can be found here. A collection of all of these reports and other materials relating to Return to Flight for the Space Shuttle fleet can be found here.

Today’s CEO (crew earth observation) photo targets, in the current LVLH attitude no longer limited by flight rule constraints on the use of the Lab nadir/science window, except for the shutter closure and condensation-prevention plan (limited to 90 min. in 24 hours), were Dust event, Namibia, SW Africa (Dynamic event.  Strong Santa Ana-type winds are blowing offshore in Namibia, driving dust plumes out to sea [visible faintly on satellite imagery].  These warm winter winds give the Namib coast its hottest temperatures of the year in winter.  The dust plumes may persist), Nyiragongo Volcano, Democratic Rep Congo (Dynamic event.  Nyiragongo started erupting ash and steam on 7/12/04, and has been active intermittently for all of 2004.  In January 2002 lava from the volcano flooded the main street of the lakeside town of Goma at the foot of the volcano.  Looking a touch left, due north of Lake Kivu), Pilcomayo River dynamics, N Argentina (good pass across the E [downstream] end of the world’s largest megafan/inland delta.  A mapping pass just left of track will reveal any changes of river course in this highly dynamic region.  Sixty percent overlap of images will allow stereoscopic viewing), and High Central Andean Glaciers (crew was to shoot any ice-capped volcanoes in sight.  Ice caps on several high glaciers have been cored and reveal a climatic record of changing snow precipitation and dust deposition for 25,000 years).

CEO images can be viewed at these websites:

See also the website “Space Station Challenge” at:

To view the latest photos taken by the expedition 9 crew visit:

U.S. & Russian Segment Status  (as of today, 1:00pm EDT)

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):

  • Elektron O2 generator is On (16A, =lowest setting).  Vozdukh CO2 scrubber is On.  U.S. CDRA CO2 scrubber is Off.  TCCS (trace contaminant control subsystem) is operating.  SM Gas Analyzer has been calibrated and is used for ppO2 & ppCO2 monitoring.  MCA (major constituents analyzer) is in Life Extending Mode (LEM).  BMP Harmful Impurities unit: absorbent bed #1 in Purify mode, bed #2 in Purify mode.  RS air conditioner SKV-1 is Off, SKV-2 is Off (SM panel mods completed 4/8; SKV-2 activation failed 4/20; is still considered failed).  SFOG slot #2 fan suspect (not usable).
  • SM Working Compartment:  Pressure (mmHg) — 752; temperature (deg C) — 23.1; ppO2 (mmHg) — 166.6; ppCO2 (mmHg) — 4.5.
  • SM Transfer Compartment:  Pressure (mmHg) — 751; temperature (deg C) — 19.1.
  • FGB Cabin:  Pressure (mmHg) — 756; temperature (deg C) — 21.3.
  • Node:  Pressure (mmHg) — 748.3; temperature (deg C) — 22.3 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • U.S. Lab:  Pressure (mmHg) — 750.4; temperature (deg C) — 24.9; ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • Joint Airlock (Equip. Lock):   Pressure (mmHg) — 750.7; temperature (deg C) — 25.1; shell heater temp (deg C) — n/a, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • (n/a = data not available)

Electrical Power Systems (EPS):

  • Both P6 channels fully operational.  BGA (beta gimbal assembly) 2B and 4B both in directed position (Dual angle/”blind” mode, non solar-tracking, biased for drag reduction).
  • SM batteries:  Battery #6 is in “Cycle” mode; all other batteries (7) are on line in “Partial Charge” mode.  
  • FGB batteries:  Battery #4 is off line; all other batteries (5) are on line in “Partial Charge” mode.
  • Plasma Contactor Unit PCU-1 is in Standby mode; PCU-2 is in Standby mode.

Command & Data Handling Systems:

  • C&C-1 MDM is prime, C&C-2 is backup, and C&C-3 is in standby.
  • GNC-2 MDM is prime; GNC-1 is backup.
  • INT-2 is operating; INT-1 is Off.
  • EXT-1 is On (primary), EXT-2 is Off (backup).
  • LA-1, LA-2 and LA-3 MDMs are all operating.
  • PL-1 MDM is Off; PL-2 MDM is Operational.
  • APS-1 (automated payload switch #1) and APS-2 are both On.
  • SM Terminal Computer (TVM): 3 redundant lanes (of 3) operational.
  • SM Central Computer (TsVM): 3 redundant lanes (of 3) operational.
  • FGB MDM-1 is Off (failed, 11/21/03); MDM-2 is Operational.

Propulsion System:

  • Total propellant load available: 3928 kg (8660 lb) as of 7/1/04;  [SM(552) + FGB(2772) + Progress M(639)].  (Capabilities: SM — 860 kg; FGB — 6120 kg).

Attitude Control Systems:

  • 3 CMGs on-line (CMG-1 failed, since 6/6/02; CMG-2’s RPC-17 failed 4/21/04; was replaced 6/30/04).
  • State vector source — U.S. SIGI-1 (GPS)
  • Attitude source — U.S. SIGI-1 (GPS)
  • Angular rate source — RGA-1

Flight Attitude:

  • LVLH XVV (local vertical/local horizontal = “earth-fixed”: z-axis in local vertical, x-axis in velocity vector [yaw: -10 deg, pitch: -9 deg, roll: 0 deg]), with CMG/TA (thruster assist) Momentum Management, until 6/28, following the EVA.

Communications & Tracking Systems:

  • FGB MDM-1 is powered Off; FGB MDM-2 is operational.
  • All other Russian communications & tracking systems are nominal.
  • S-band is operating nominally (on string 2).
  • Ku-band is operating nominally (may require a mask).
  • Audio subsystem is operating nominally (IAC-1 is prime, IAC-2 is off).
  • Video subsystem operating nominally.
  • HCOR (high-rate communications outage recorder) is operating nominally.


  • SSRMS/Canadarm2 based at Lab PDGF/LEE A, operational on redundant string, off on prime.
  • MBS: KA (keep alive) power on both strings. 
  • MT: latched and mated at WS4. 
  • POA: KA power on both strings.
  • RWS (robotics workstations): Lab RWS is On (DCP connected); Cupola RWS is Off.

ISS Location NOW

Full Size/Update
Real Time ISS TrackerMore Links

ISS Orbit  (as of this morning, 6:23am EDT [= epoch]):

  • Mean altitude — 359.6 km
  • Apogee height — 363.8 km
  • Perigee height — 355.5 km
  • Period — 91.7 min.
  • Inclination (to Equator) —  51.6302 deg
  • Eccentricity — 0.0006174
  • Solar Beta Angle — 64.3 deg
  • Orbits per 24-hr. day — 15.70
  • Mean altitude loss in last 24 hours — 75 m
  • Revolutions since FGB/Zarya launch (Nov. ’98)  — 32275

ISS Altitude History

Apogee height Mean AltitudePerigee height

ISS Altitude History

For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see In addition, information on International Space Station sighting opportunities can be found at on NASA’s Human Spaceflight website. The current location of the International Space Station can be found at at NASA’s Marshall Space Flight Center. Additional satellite tracking resources can be found at

SpaceRef staff editor.